# Differential Privacy for Deep Learning in Medicine

### v1

@@ 7.1 Federated and Mechanism-Agnostic Privacy Approaches
Table 5: Summary of studies applying alternative DP mechanisms beyond DP-SGD. Each entry includes the data modality, downstream task, and the DP mechanism and accounting method used. These include applications of local DP, encryption-based aggregation, federated learning (FL) variants, and generative modeling frameworks across a range of clinical modalities and tasks. AC: advanced composition; AD: Alzheimer’s disease; EHR: electronic health record; EM: exponential mechanism; GDP: Gaussian differential privacy; GM: Gaussian mechanism; PD: Parkinson’s disease; NA: information not available; LC: linear composition; LM: Laplace mechanism; MA: moments accountant; PLD: privacy loss distribution; RDP: Rényi DP accountant; RR: randomized response.
Several studies extended DP to FL settings, where collaborative training is performed without centralizing patient data. Alsenani et al. [93] introduced the FAItH framework, which combines local and centralized DP with secure multiparty computation for sensor data clustering. Gong et al. [95] used Gaussian noise with moment accounting in a federated intensive care unit monitoring system. Yuan et al. [97] employed Gaussian DP in a federated chest X-ray classification pipeline. Riess et al. [39] adapted federated training for MRI classification using complex-valued neural networks.
Some studies (e.g., [100, 115]) demonstrated FL-based medical models under DP constraints but did not report full accounting details. This lack of standardization highlights ongoing challenges in reproducibility and transparency.
Outside of FL, many studies used alternative mechanisms in centralized or generative contexts. Laplace mechanisms were used in tasks such as pathological speech classification [126], cancer risk modeling [129], and drug sensitivity prediction [123]. Exponential mechanisms were applied in EHR and clinical text classification tasks [131, 92], often in combination with Laplace noise or local perturbation.
Some studies, such as Wu et al. [135], adopted one-shot local DP strategies where perturbation occurs per sample, requiring no composition accounting. Similarly, Gwon et al. [96] used randomized response mechanisms for synthetic EHR generation, emphasizing privacy at the point of data release rather than during model training.
Generative modeling also featured prominently. He et al. [121] and Zhang et al. [136] synthesized RNA-Seq and EHR data under DP. Notably, Sun et al. [98] applied noise to the loss gradients—rather than individual per-sample gradients—when training a generator, a setup inspired by the DP-GAN framework. This contrasts with DP-SGD-based GANs (e.g., [82]) that apply DP to per-sample gradients in the discriminator.
Other applications included DP for data sharing [136], motif discovery in DNA sequences [116], and survival analysis using exponential and Laplace mechanisms [120, 133].
Despite their diversity, alternative mechanisms often suffer from inconsistent reporting. Many studies omit essential parameters, such as the privacy budget $(\epsilon)$, the delta value, or the type of accountant used. Mechanisms like randomized response or one-step perturbation—while avoiding the need for composition—still require careful tuning to balance utility and privacy. Encryption-based designs introduce additional complexity through communication overhead and synchronization requirements.
Nevertheless, these approaches highlight a shift toward more flexible, mechanism-aware privacy strategies. By leveraging differentially private techniques suited to specific tasks and data types, these methods offer scalable alternatives to DP-SGD—particularly in federated, generative, or low-resource settings. Moving forward, clearer guidelines for reporting, accounting, and mechanism selection will be essential for advancing fair, privacy-preserving AI in healthcare.
7.2 Defenses Against Attacks
@@
Almadhoun et al. [114] GP Attribute inference attack, MIA DP-SGD
Almadhoun et al. [114] GP MIA DP-SGD
Bingzhe et al. [77] GP Model-inversion attack DP-SGD
Díaz et al. [117] Weight perturbation Client inference attack Beyond DP-SGD
DPFL [80] GP MIA DP-SGD
Fan et al. [66] GP MIA DP-SGD
Gwon et al. [96] Input perturbation MIA Beyond DP-SGD
Hatamizadeh et al. [69] GP GIA (BN-aware inversion) DP-SGD
Hyunwook et al. [137] Model-based perturbation Re-identification risk assessment Beyond DP-SGD
Kaissis et al. [4] GP Model-inversion attack DP-SGD
Liangrui et al. [78] GP GIA DP-SGD
Riess et al. [39] GP GIA DP-SGD
Shunrong et al. [65] GP MIA DP-SGD
Sun et al. [98] GP Identity disclosure attack DP-SGD
Sun et al. [98] GP Attribute disclosure attack DP-SGD
Tramèr et al. [132] Output perturbation MIA (theoretical) Beyond DP-SGD
Veeraragavan et al. [133] GP MIA DP-SGD
You et al. [93] Output perturbation MIA Beyond DP-SGD
Ziller et al. [111] GP Reconstruction attack DP-SGD
@@
Membership inference attacks (MIAs) attempt to infer whether a specific sample was part of the model’s training dataset—a particularly serious risk in healthcare, where inclusion in a disease cohort may itself reveal sensitive clinical information. Multiple studies applied DP-SGD to protect against MIAs in classification tasks using imaging, sensor, and tabular data [80, 66, 65, 133]. These defenses typically introduced Gaussian noise into gradient updates, reducing attack success rates even at moderate $\epsilon$ values.
However, the strength of protection varied across settings. Fan et al. [66] noted that class imbalance and shallow architectures increased MIA vulnerability under DP, emphasizing the need to consider model and data characteristics when tuning privacy parameters. Almadhoun et al. [114] confirmed that even with DP-SGD, certain MIAs could still succeed under weak noise or limited clipping, highlighting the importance of robust DP configurations.
@@
Beyond MIAs, several studies evaluated inversion-based threats, where attackers attempt to reconstruct sensitive inputs—such as medical images or EHR records—by analyzing shared gradients. These attacks are especially relevant in federated learning, where gradient exchange is common. For example, Hatamizadeh et al. [69] and Liangrui et al. [78] demonstrated that DP-SGD defenses could substantially degrade inversion quality but not eliminate leakage entirely. Ziller et al. [111] observed similar results in reconstruction attacks on multi-institutional imaging models.
Alternative perturbation schemes, such as weight noise [117] or complex-valued gradients [39], were also explored as mitigation strategies. Still, these methods remain under-tested and lack standardized evaluation criteria, especially under stronger adversary assumptions (e.g., white-box gradient access).
@@
LDP, which perturbs data before model access, was investigated in a few studies (e.g., [93, 96, 132]). While theoretically strong, LDP was shown to be insufficient in practice against sophisticated attacks, including attribute inference and partial MIAs. These findings underscore that DP defenses must often extend beyond input perturbation, especially in distributed learning settings.
@@
A recurring recommendation across studies is the use of hybrid defenses—combining DP with cryptographic methods like secure aggregation or multiparty computation [93, 135]. These layered strategies offer stronger guarantees but introduce practical constraints, including communication overhead and increased system complexity.
Despite growing interest in attack resilience, several gaps remain. First, the majority of studies focus on a narrow set of attacks—most commonly MIAs—while overlooking others like training-time poisoning or client-level inference in federated setups. Second, many reports lack standardized threat models or omit key experimental details (e.g., batch size, gradient access scope), limiting reproducibility. Third, defenses are rarely evaluated under varying $\epsilon$ levels, making it difficult to assess tradeoffs between privacy strength and robustness.
DP offers a foundational layer of protection against common inference attacks in medical DL. However, achieving effective defense requires not only tuning $\epsilon$, but also understanding how model architecture, data distribution, and perturbation method interact with adversarial goals. Future work should prioritize comprehensive, reproducible evaluations that benchmark attack resilience under clinically realistic settings and privacy constraints.
8 Discussion
@@
Despite the theoretical strengths of DP [10], practical implementations face well-documented tradeoffs. Performance consistently degrades under stricter privacy budgets, particularly at $\epsilon\approx 1$, where noise injection frequently leads to convergence failure or sharp accuracy declines [5, 68, 81]. However, several studies demonstrated that with strong pretraining [5, 74], stable architectures [13, 5, 71], and careful normalization [81], models can maintain clinically acceptable utility even under strong privacy guarantees (ϵitalic-$\epsilon\approx$10). Imaging tasks, particularly chest X-ray and CT classification, showed resilience to DP noise [13, 5], while domains such as histopathology [58], dermoscopy [68], and pathological speech [81] proved more vulnerable, largely due to smaller datasets or higher input variability.
Fairness remains underexplored relative to utility. A handful of studies—most notably those by Tayebi Arasteh et al. [13, 5, 81]—conducted thorough evaluations of subgroup disparities under varying privacy levels, revealing that DP can exacerbate demographic gaps unless explicitly mitigated. However, fairness is rarely treated as a core training objective. Intersectional effects are largely absent, and fairness in regression or survival tasks remains rarely assessed [66, 133]. Fairness-aware optimization methods such as reweighting or constraint-based training are virtually unexplored in the current literature. Metrics like subgroup AUC, underdiagnosis rate, and statistical parity difference have been used to track disparities [13, 71, 61], but no consensus exists on which are most clinically meaningful or how to interpret them under noisy predictions.
Alternative privacy mechanisms beyond DP-SGD—including Laplace and exponential noise [126, 92, 120], federated aggregation [93, 95, 97], and local perturbation [96, 135]—offer increased flexibility and often better alignment with deployment constraints. Generative models, such as GANs and diffusion models, were applied to EHR and medical image synthesis under DP constraints [121, 119, 98]. However, many studies omit essential parameters like $(\epsilon,\delta)$ or the privacy accountant used, complicating reproducibility and comparison.
@@
We also reviewed the threat landscape for privacy-preserving models. Empirical evaluations show that DP can significantly mitigate attacks such as membership inference and gradient inversion [80, 66, 69, 111], but protections are highly context-dependent. Fan et al. [66] emphasized that class imbalance and shallow architectures increased vulnerability. Almadhoun et al. [114] demonstrated residual attack success under weak noise, even with DP-SGD. Some studies also evaluated reconstruction attacks in FL [78, 138], identity and attribute disclosure risks [98], and defenses using weight perturbation [117] or complex-valued networks [39]. Importantly, LDP alone was shown to be insufficient under adversarial conditions [96, 132].
@@
This review has several limitations. While our search covered major venues and was supplemented by manual inclusion of key preprints, the rapid pace of publication means that recent work may have been missed. In addition, while we systematically included all eligible studies applying DP to medical DL, some clinical areas—such as genomics [63, 130] or survival modeling [66, 133]—are still underrepresented in the literature itself. Finally, our review focused primarily on technical methods and performance outcomes; perspectives from clinicians, patients, and data custodians were outside the scope but remain essential for translating privacy-preserving models into practice.
Several future directions emerge. First, fairness evaluations should become standard practice in DP studies, ideally spanning multiple $\epsilon$ levels and well-justified subgroup definitions. Second, benchmark datasets with demographic annotations and support for intersectional analysis would facilitate robust, reproducible fairness evaluations. Third, new methods that integrate privacy and fairness directly into training—such as group-aware DP optimizers or adaptive clipping—could help ensure that protections do not exacerbate health inequities. Fourth, clearer documentation of DP parameters, accountant types, and performance tradeoffs will improve transparency and support rigorous benchmarking.
Differential privacy offers a robust framework for protecting individual data in clinical AI models. However, privacy alone does not guarantee equitable or reliable deployment. A shift toward integrated design—where privacy, fairness, and utility are optimized together—is essential for developing trustworthy, real-world-ready medical AI. This review provides a foundation for that goal, clarifying what is already known and identifying what remains to be addressed.
Additional information

### v2

@@ Results
Table 5: Summary of studies applying alternative DP mechanisms beyond DP-SGD. Each entry includes the data modality, downstream task, and the DP mechanism and accounting method used. These include applications of local DP, encryption-based aggregation, federated learning (FL) variants, and generative modeling frameworks across a range of clinical modalities and tasks. AC: advanced composition; AD: Alzheimer’s disease; EHR: electronic health record; EM: exponential mechanism; GDP: Gaussian differential privacy; GM: Gaussian mechanism; PD: Parkinson’s disease; NA: information not available; LC: linear composition; LM: Laplace mechanism; MA: moments accountant; PLD: privacy loss distribution; RDP: Rényi DP accountant; $R$: randomized response.
The first pattern replaces gradient perturbation with privacy-preserving knowledge transfer. In the private aggregation of teacher ensembles (PATE) framework [PATE], sensitive training data are partitioned across disjoint teacher models,
$D=\bigcup_{k=1}^{K}D_{k},\qquad D_{i}\cap D_{j}=\emptyset\text{ for }i\neq j$ (24)
and each teacher produces a class prediction. Votes for each class are tallied,
$v_{c}(x)=\sum_{k=1}^{K}\mathbb{I}\big[T_{k}(x)=c\big],$ (25)
after which Laplace noise is added to the vote counts,
$\tilde{v}_{c}(x)=v_{c}(x)+\mathrm{Lap}(1/\lambda),$ (26)
and the final label exposed to the student model is the noisy majority vote,
$\hat{y}(x)=\arg\max_{c}\tilde{v}_{c}(x).$ (27)
By training only on noisy labels, student models never interact with private data. Studies using PATE, particularly for GAN-based synthetic data [PATEvsDPSGDGAN], reported better utility at small $\epsilon$ than DP-SGD when data could be partitioned.
A second group of studies applied DP in federated or decentralized settings, enabling collaborative training without sharing data. DP noise was injected either locally at each client or centrally after aggregation. Applications included federated clustering of sensor data [23FAItH], intensive care unit (ICU) monitoring [27gong2023federated], pneumonia classification from chest X-rays [49yuan2019collaborative], and MRI models using complex-valued networks [73Complex-ValuedFederatedLearningwithDifferentialPrivacy]. However, some papers (e.g., [60DifferentialPrivateFederatedTransferLearningforMentalHealthMonitoringinEverydaySettings, 9Dp-FedAvg]) did not report $(\epsilon,\delta)$ or the privacy accountant, limiting reproducibility.
A third pattern perturbed data or outputs directly rather than modifying gradients. Laplace and exponential mechanisms were used for speech classification [11ADDETECTOR], EHR and clinical text modeling [17NLPICD-10, 31AZIZ2022104113], and survival prediction [61fedeli2021privacy]. Local differential privacy was applied in tabular and EHR sharing [79wu2021privacy, 42GWON2024107738], avoiding composition accounting but typically reducing accuracy.
A fourth direction generated DP-synthetic data, decoupling downstream model training from private records. GANs and diffusion models were trained under DP constraints for RNA-sequence [12DP-SL-GAN], tabular EHR [78zhang2022differential], and CTGAN-based tabular data [83DP-CTGAN]. Sun et al. [39SUN2023104404] added noise to generator loss gradients rather than per-sample gradients, improving stability at low privacy budgets.
Across these studies, we observed substantial heterogeneity in reporting: more than one-third omitted essential details such as $(\epsilon,\delta)$ or the accountant used. Hybrid designs combining encryption and DP [DPMPC, DPplusMPC] illustrate that DP is increasingly implemented as a system-level deployment strategy, not only a gradient-level mechanism.
Defenses against attacks
@@
Almadhoun et al. [76Inferenceattacksagainstdifferentiallyprivatequery] GP Attribute inference attack, MIA DP-SGD
Almadhoun et al. [76Inferenceattacksagainstdifferentiallyprivatequery] GP MIA DP-SGD
Bingzhe et al. [64wu2019p3sgd] GP Model-inversion attack DP-SGD
Díaz et al. [85diaz2025metric] Weight perturbation Client inference attack Beyond DP-SGD
DPFL [3DPFRCovid19] GP MIA DP-SGD
Fan et al. [26fan2023mitigating] GP MIA DP-SGD
Gwon et al. [42GWON2024107738] Input perturbation MIA Beyond DP-SGD
Hatamizadeh et al. [33hatamizadeh2023gradient] GP GIA DP-SGD
Hyunwook et al. [45kim2024synthetic] Model-based perturbation Re-identification risk assessment Beyond DP-SGD
Kaissis et al. [66kaissis2021end] GP Model-inversion attack DP-SGD
Liangrui et al. [63FedDP] GP GIA DP-SGD
Riess et al. [73Complex-ValuedFederatedLearningwithDifferentialPrivacy] GP GIA DP-SGD
Shunrong et al. [50DPFedSAM-Meas] GP MIA DP-SGD
Sun et al. [39SUN2023104404] GP Identity disclosure attack DP-SGD
Sun et al. [39SUN2023104404] GP Attribute disclosure attack DP-SGD
Tramèr et al. [80DifferentialPrivacywithBoundedPriors] Output perturbation MIA Beyond DP-SGD
Veeraragavan et al. [81raghavan2024differentially] GP MIA DP-SGD
You et al. [23FAItH] Output perturbation MIA Beyond DP-SGD
Ziller et al. [67ziller2024reconciling] GP Reconstruction attack DP-SGD
@@
Membership inference attacks (MIAs) attempt to infer whether a specific sample was part of the model’s training dataset, a particularly serious risk in healthcare, where inclusion in a disease cohort may itself reveal sensitive clinical information. Multiple studies applied DP-SGD to protect against MIAs in classification tasks using imaging, sensor, and tabular data [3DPFRCovid19, 26fan2023mitigating, 50DPFedSAM-Meas, 81raghavan2024differentially]. These defenses typically introduced Gaussian noise into gradient updates, reducing attack success rates even at moderate $\epsilon$ values. However, the strength of protection varied across settings. Fan et al. [26fan2023mitigating] noted that class imbalance and shallow architectures increased MIA vulnerability under DP, emphasizing the need to consider model and data characteristics when tuning privacy parameters. Almadhoun et al. [76Inferenceattacksagainstdifferentiallyprivatequery] confirmed that even with DP-SGD, certain MIAs could still succeed under weak noise or limited clipping, highlighting the importance of robust DP configurations.
@@
Beyond MIAs, several studies evaluated inversion-based threats, where attackers attempt to reconstruct sensitive inputs, such as medical images or EHR records, by analyzing shared gradients. These attacks are especially relevant in federated learning, where gradient exchange is common. For example, Hatamizadeh et al. [33hatamizadeh2023gradient] and Liangrui et al. [63FedDP] demonstrated that DP-SGD defenses could substantially degrade inversion quality but not eliminate leakage entirely. Ziller et al. [67ziller2024reconciling] observed similar results in reconstruction attacks on multi-institutional imaging models. Alternative perturbation schemes, such as weight noise [85diaz2025metric] or complex-valued gradients [73Complex-ValuedFederatedLearningwithDifferentialPrivacy], were also explored as mitigation strategies. Still, these methods remain under-tested and lack standardized evaluation criteria, especially under stronger adversary assumptions (e.g., white-box gradient access).
@@
LDP, which perturbs data before model access, was investigated in a few studies (e.g., [23FAItH, 42GWON2024107738, 80DifferentialPrivacywithBoundedPriors]). While theoretically strong, LDP was shown to be insufficient in practice against sophisticated attacks, including attribute inference and partial MIAs. These findings underscore that DP defenses must often extend beyond input perturbation, especially in distributed learning settings.
@@
Finally, a recurring recommendation across studies is the use of hybrid defenses, combining DP with cryptographic methods like secure aggregation or multiparty computation [23FAItH, 79wu2021privacy, MPC, SMPC, DPMPC, DPplusMPC]. These layered strategies offer stronger guarantees but introduce practical constraints, including communication overhead and increased system complexity.
Discussion
@@
Despite the theoretical strengths of DP [dwork2014algorithmic], practical implementations face well-documented tradeoffs. Performance consistently degrades under stricter privacy budgets, particularly at $\epsilon\approx 1$, where noise injection frequently leads to convergence failure or sharp accuracy declines [22tayebi2024preserving, 53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning, 72arasteh2024differentialprivacyenablesfair]. However, several studies demonstrated that with strong pretraining [22tayebi2024preserving, 47mehmood2024balancing], stable architectures [13TayebiDomainTransfer, 22tayebi2024preserving, 88FairandPrivateCTContrastAgentDetection], and careful normalization [72arasteh2024differentialprivacyenablesfair], models can maintain clinically acceptable utility even under strong privacy guarantees ($\epsilon\approx$10). Imaging tasks, particularly chest X-ray and CT classification, showed resilience to DP noise [13TayebiDomainTransfer, 22tayebi2024preserving], while domains such as histopathology [36adnan2022federated], dermoscopy [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning], and pathological speech [72arasteh2024differentialprivacyenablesfair] proved more vulnerable, largely due to smaller datasets or higher input variability.
Beyond performance considerations, our findings highlight that DP introduces constraints that persist after training and influence real-world deployment [cormode2018privacy, busch2025privacy, 66kaissis2021end]. Once a model is trained with DP, privacy guarantees remain tied to how often the model, or the underlying dataset, is accessed. Multiple users querying a DP-trained model, or repeated inference on the same data (e.g., auditing, calibration, federated reuse), incrementally consume the privacy budget through composition of privacy loss [dwork2014algorithmic]. In adversarial scenarios, a malicious user could issue strategically engineered repeated queries to partially cancel injected noise or accelerate privacy depletion. As a result, DP must be treated as a finite, depletable resource over the model’s lifecycle rather than a one-time property of training [66kaissis2021end, 67ziller2024reconciling]. Operational deployments therefore require governance mechanisms, such as per-user privacy budgeting, access authentication, query logging, audit trails, and automatic rejection of further queries once the remaining privacy budget is exhausted [cormode2018privacy, abowd2018us, near2025guidelines]. These constraints position DP not only as a mathematical guarantee but as a system-level design requirement that influences how DP models are shared, reused, and exposed in hospital or federated environments.
The privacy parameters $\epsilon$ and $\delta$ determine the strength of a DP guarantee, yet their practical implications are often under‐reported. A smaller $\epsilon$ corresponds to stronger privacy, meaning the model’s outputs are less sensitive to any single training example, but achieving low $\epsilon$ requires adding more noise, which reduces model performance. Across the reviewed studies, strict budgets ($\epsilon\approx 1$) frequently led to convergence failure or sharp utility drops, whereas moderate budgets ($\epsilon\approx 10$) preserved clinical accuracy in structured imaging tasks. The $\delta$ parameter, typically set to a negligible value (e.g., $1/N$ where $N$ is dataset size), represents the probability that the privacy guarantee may not hold; however, many papers report $\epsilon$ without disclosing $\delta$ or the privacy accountant used. These observations show that meaningful interpretation of privacy-utility tradeoffs requires reporting both parameters and contextualizing them in terms of clinical task fidelity rather than treating $\epsilon$ in isolation.
Fairness remains underexplored relative to utility. Only a subset of studies explicitly audit subgroup disparities under DP [13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair, 88FairandPrivateCTContrastAgentDetection, 74chin2020generation]. These works focus mainly on classification and binary protected attributes (e.g., age, sex, race). Noise introduced by DP can interact with data imbalance and amplify disparities, particularly for underrepresented groups or tasks already exhibiting poor performance [cummings2019compatibility, tran2021decision]. For example, Tayebi Arasteh et al. showed that DP disproportionately degraded performance for certain age groups in speech classification [72arasteh2024differentialprivacyenablesfair]. However, a recent multi-institutional evaluation found that performance degradation under DP was uncorrelated with subgroup size and instead depended on task difficulty, with clinically complex or rare conditions affected most; at moderate privacy levels ($\epsilon\approx 5$–10), fairness remained stable, and some underrepresented groups even benefited from stronger privacy guarantees [22tayebi2024preserving]. These results echo broader fairness literature showing that inequity often stems from dataset and labeling structure [Mehrabi2021Survey, obermeyer2019dissecting]. Despite these insights, fairness is rarely a design objective. Nearly all studies report disparities only post-hoc and at a single $\epsilon$ value, limiting understanding of privacy–fairness tradeoffs. Intersectional fairness (e.g., older female patients) and non-classification tasks such as regression and survival prediction remain largely unexamined [26fan2023mitigating, 81raghavan2024differentially]. Fairness-aware optimization strategies, such as reweighting, constraint-based objectives, or adversarial debiasing, are virtually absent, despite maturity in the broader AI fairness community [Mehrabi2021Survey]. Metrics such as subgroup AUC, underdiagnosis rate, and statistical parity difference have been used [13TayebiDomainTransfer, 88FairandPrivateCTContrastAgentDetection, 74chin2020generation], but no consensus exists on which metrics are most clinically meaningful or how to interpret them under noisy predictions. To prevent fairness regressions, future DP pipelines should report subgroup performance across multiple $\epsilon$ levels, justify protected attribute selection based on clinical relevance, monitor fairness during deployment, and treat fairness as a first-class objective rather than an after-the-fact audit [Mehrabi2021Survey, obermeyer2019dissecting].
@@
We also reviewed the threat landscape for privacy-preserving models. Empirical evaluations show that DP can substantially mitigate attacks such as membership inference and gradient inversion [3DPFRCovid19, 26fan2023mitigating, 33hatamizadeh2023gradient, 67ziller2024reconciling], but protections are highly context-dependent. Fan et al. [26fan2023mitigating] emphasized that class imbalance and shallow architectures increased vulnerability. Almadhoun et al. [76Inferenceattacksagainstdifferentiallyprivatequery] demonstrated residual attack success under weak noise, even with DP-SGD. Some studies also evaluated reconstruction attacks in FL [63FedDP, truhn2024encrypted], identity and attribute disclosure risks [39SUN2023104404], and defenses using weight perturbation [85diaz2025metric] or complex-valued networks [73Complex-ValuedFederatedLearningwithDifferentialPrivacy]. Importantly, LDP alone was shown to be insufficient under adversarial conditions [42GWON2024107738, 80DifferentialPrivacywithBoundedPriors]. Despite growing interest in attack resilience, several gaps remain. First, the majority of studies focus on a narrow set of attacks, most commonly MIAs, while overlooking others like training-time poisoning or client-level inference in federated setups. Second, many reports lack standardized threat models or omit key experimental details (e.g., batch size, gradient access scope), limiting reproducibility. Third, defenses are rarely evaluated under varying $\epsilon$ levels, making it difficult to assess tradeoffs between privacy strength and robustness.
@@
This review has several limitations. First, while our search covered major venues and was supplemented by manual inclusion of key preprints, the rapid pace of publication means that very recent studies may not have been captured. Second, although we systematically included all eligible work applying DP to medical DL, the evidence base remains uneven: some clinical areas, such as genomics [25zhang2022secure, 68Representationtransferfordifferentiallyprivatedrugsensitivityprediction] and survival modeling [26fan2023mitigating, 81raghavan2024differentially], are still underrepresented in the literature. Third, the overwhelming majority of empirical studies use DP-SGD; therefore, most quantitative synthesis naturally focuses on this mechanism, while alternative approaches (e.g., local DP, PATE, DP-synthetic data generation) are summarized conceptually due to limited and heterogeneous reporting. This reflects a limitation of the current field rather than a methodological bias of the review [de2022unlocking, chua2024private]. Finally, our synthesis centers on technical and performance dimensions; perspectives from clinicians, patients, and data custodians were outside the scope but remain essential for translating privacy-preserving models into practice.
Several priorities emerge from this synthesis. First, given the systematic fairness gaps observed across studies, future DP work should include fairness evaluation across multiple $\epsilon$ levels [Mehrabi2021Survey, obermeyer2019dissecting]. Second, benchmark datasets with demographic annotations and support for intersectional subgrouping are needed to enable comparable fairness audits under DP constraints [mitchell2019model, gebru2021datasheets]. Third, future DP methods should integrate fairness during training rather than treating it as an after-the-fact audit; emerging techniques such as group-aware clipping or adaptive noise allocation offer promising directions [cummings2019compatibility, tran2021decision]. Fourth, reproducibility requires clearer reporting: studies should disclose $(\epsilon,\delta)$, the accountant type, clipping norm, and performance at multiple privacy levels. Without such reporting, privacy–utility tradeoffs cannot be interpreted or compared [dwork2014algorithmic, opacus_neurips2021]. DP offers a principled mechanism to limit individual information leakage, but privacy alone does not guarantee equitable or reliable deployment. The evidence reviewed here shows that model utility, fairness, and attack resilience depend not only on privacy budgets but also on architectural choices, pretraining, dataset structure, and evaluation practices. Moving forward, integrated design, optimizing privacy, fairness, and utility together rather than treating them as isolated objectives, will be necessary for trustworthy clinical AI [tabassi2023artificial, act2024eu].
Based on the evidence reviewed, deploying DP in clinical AI requires coordinated technical and governance decisions. For developers, pretraining and compact CNN architectures (e.g., ResNet-9/18) repeatedly appeared in studies that maintained good utility under DP. Several works used group normalization to avoid batch-statistics leakage, though current evidence does not establish a universal best practice across modalities. Fairness auditing should become standard, with subgroup reporting across multiple privacy budgets and justification of protected attributes. For healthcare institutions, privacy budget should be treated as a depletable resource after deployment, requiring governance controls such as authenticated access, per-user query limits, logging, and audit trails. For policymakers and regulators, transparent reporting of $(\epsilon,\delta)$, the accountant used, clipping norm, and performance across privacy levels should be required to ensure comparability across systems [66kaissis2021end, cummings2018role, near2025guidelines]. These practices move DP from a theoretical guarantee to an operational framework for safe, reproducible, and equitable deployment in clinical environments.
Methods
This scoping review was conducted in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [tricco2018prisma, page2021prisma] to ensure methodological rigor and transparency [buess2025large]. The complete PRISMA-ScR checklist is provided in Supplementary Table 1. The methodological approach to data collection, article selection, and results synthesis is described in detail below.
Information sources
We systematically searched four electronic databases: PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. Database queries were performed on March 1, 2025, and search results were imported into Rayyan software [ouzzani2016rayyan], a tool designed to facilitate the systematic screening of literature and duplicate removal. A detailed description of the database-specific search queries is provided in Supplementary Table 2.
Search strategy
The search strategy was carefully designed to identify publications at the intersection of DP, DL, and medical applications. Specifically, the search strings combined three primary keyword groups: (i) differential privacy techniques (e.g., “differential privacy,” “differentially private,” “DP-SGD”), (ii) deep learning and related machine learning methodologies (e.g., “deep learning,” “machine learning,” “neural network,” “artificial intelligence”), and (iii) healthcare and clinical applications (e.g., “medical,” “clinical,” “healthcare,” “biomedical,” “EHR,” “imaging,” “radiology,” “text,” “speech”).
Inclusion and exclusion criteria
The literature selection involved multiple screening steps: (i) Initial screening included automated duplicate removal and manual title and abstract screening in Rayyan [ouzzani2016rayyan]. Articles were excluded if they lacked relevance to DP or healthcare applications. (ii) Full-text reviews were conducted on the remaining articles to ensure alignment with inclusion criteria, specifically the empirical application of DP with DL techniques in medical contexts. Articles with purely theoretical content, lacking empirical evaluation, or applied to non-healthcare domains were excluded. The selection process ensured proportional representation across different medical application domains, including medical imaging, electronic health records, predictive analytics, and personalized medicine.
Data availability
No datasets were generated or analyzed during the current study. This is a scoping review based entirely on previously published literature.
Code availability
No custom code was developed or used in this study. All analyses were conducted through manual review of existing publications.
Acknowledgements
We acknowledge financial support by Deutsche Forschungsgemeinschaft. DT was supported by grants from the DFG (NE 2136/3-1, LI3893/6-1, TR 1700/7-1) and is supported by the German Federal Ministry of Education (TRANSFORM LIVER, 031L0312A; SWAG, 01KD2215B) and the European Union’s Horizon Europe and innovation programme (ODELIA [Open Consortium for Decentralized Medical Artificial Intelligence], 101057091). The funders played no role in the design or execution of the study. The authors of this work take full responsibility for its content.
Author contributions